In a recent poll, 140 people were asked if they liked dogs, and 22% said they did. Find the Margin of Error for this poll, at the 95% confidence level. Give your answer to four decimal places if possible.
Answer #1

Given that: n=140 , p=0.22

alpha =0.05 , From z table Zcrit= Z(alpha/2)=1.96

Margin of Error= Zcrit * \sqrt{p(1-p)/n} =1.96*\sqrt{(0.22*0.78)/140} =0.0686

ANSWER MARGIN OF ERROR=0.0686